What is the magnitude of the field if a charge of 10 C travelling at a speed of 8 × 106 m/s in a direction perpendicular to a magnetic field experiences a magnetic force of 8.7 × 10-3 N?


1
Expert's answer
2022-03-22T15:35:06-0400

The magnetic force

"F=qvB\sin\theta"

Hence

"B=\frac{F}{qv\sin\theta}\\
=\frac{8.7 \times10^{-3}}{10\times 8\times 10^6\times \sin90^\circ}=1.1\times 10^{-10}\:\rm T"